OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 2x – 5y = 9?

OpenStudy (anonymous):

what the right choice

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

When lines are parallel they have the same slope. So all you need to do is put this equation in slope-intercept form y=mx+b where m is the slope of the line. Since parallel lines have the same slope, the value of m will be the slope they're asking for.

OpenStudy (anonymous):

2x-5y=9 Solve for y. Subtract 2x from both sides: -5y=-2x+9 Divide both sides by -5: y=(2/5)x-(9/5)